You are considering opening your own restaurant. To do so, you will have to quit your current job, which pays $46k per year, and cash in your life savings of $200k, which have been in a certificate of deposit paying 6% per year. You will need this $200k to purchase equipment for your restaurant operations. You estimate that you will have to spend $4k during the year to maintain the equipment so as to preserve its market value at $200k. Fortunately, you own a building suitable for the restaurant. You currently rent out this building on a month-by-month basis for $2,500 per month. You anticipate that you will spend $50k for food, $40k for extra help, and $14k for utilities and supplies during the first year of operations. There are no other costs involved in this business. We want to know how much revenue your restaurant must bring in order to have positive economic profits.

a. What is your direct cost (or explicit cost) of operating the restaurant during the first year?

b. What is your opportunity cost (or implicit cost) of operating the restaurant during the first year?
Answer #1

a. Direct cost will include the expenditure incurred in starting the business not including the opportunity foregone-

This includes in this case Equipment cost+maintainable cost+ food cost+ utility cost+help cost=200000+4000+50000+40000+14000=$308000

b.Opportunity cost includes the cost of forgoing the opportunity so in this case will include-

Cost of Existing employment+interest on deposit+rental income from building=46000+0.06*200000+2500=46000+12000+2500=$60500